OpenStudy (anonymous):

Rewrite y = x^2 + 14x + 29 in general form.. "Completing the Square."

OpenStudy (anonymous):

\[y = x^2 + 2*7*x + 49 -49 + 29\] \[y = (x +7)^2 -20\] \[y + 20 = (x+7)^2\]

OpenStudy (anonymous):

Wait in the first equation, you mean \[y = x ^{2} +7^{2} x +49 - 49 +29\] right?